6. If cos(x) = -, and xeQIV , find the exact value of each of the following. Show all work. a.sin(2x) X b.cos (६)
Answer #1

solution given that Cosla) = Ź and 2 E ON las to find the value of sinza , We know the formwa of sinax , Sim2 x = 2 Sim CoCx50, Cosa = t-(3 sona NIM = 9-( he lason 2 205246) Ž+ Et2 = (helyson = 2+3 3 26052/7) = toim cos? (2/2) = 5 6 cos(x) = TV here